OpenStudy (anonymous):

1/4x ^{2}-1/12x-1/6=0 Solve.

OpenStudy (chaise):

This is a quadratic equation. a = 1/4 b = -1/12 c = -1/6 \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[x = \frac{-3}{2}\] or \[x=1\]

OpenStudy (anonymous):

so you are saying the is x=-3/2&1?

OpenStudy (chaise):

Nope, I am saying x = -3/2 OR x = 1 If you substitute your x value as 1, OR if you substitute your x value as -3/2 you will get your equation equal to zero. When you solve quadratic equations you (pretty much) always get two answers.

OpenStudy (anonymous):

ok thanks

OpenStudy (chaise):

If you are unfamiliar with the quadratic formula you should try watching a youtube video on quadratics.